OpenStudy (anonymous):

what is the value of 7-^3

OpenStudy (michele_laino):

hint: \[\Large {7^{ - 3}} = \frac{1}{{{7^3}}} = \frac{1}{7} \times \frac{1}{7} \times \frac{1}{7} = ...?\]

OpenStudy (anonymous):

These r the options a.-21 b.-1/343 c.1/343 or d. 343

OpenStudy (lynfran):

follow @Michele_Laino hint it will lead u to the answer @Maryjaen_14

OpenStudy (dinamix):

1/343

OpenStudy (kenialove):

What's 1/7 x 1/7 x 1/7 ? @Maryjaen_14

OpenStudy (michele_laino):

hint: \[\Large {7^{ - 3}} = \frac{1}{{{7^3}}} = \frac{1}{7} \times \frac{1}{7} \times \frac{1}{7} = \frac{1}{{7 \times 7 \times 7}} = ...?\]